OpenStudy (anonymous):

4 adults and 5 children attended a funraising event. Tickets for adults were $3 more than child tickets. If the total money spent for admission at the fundraiser was $39, what was the cost of a CHILD's ticket?

OpenStudy (kamille):

hey, you need to make an equation and solve it to find the price of the ticket. Since you need to find how much did child's ticket cost, let's say it costed x $, then adult ticket costed x+3. since you know that there were 4 adults and 5 children,you can make the equation: 5x+4(x+3)=39 and find out what was the cost of the children ticket

OpenStudy (kamille):

could you please solve this equation for me, I will check the answer

OpenStudy (anonymous):

cant thts why i asked u guys

OpenStudy (kamille):

\[5x+4x+12=39\] \[9x=39-12\] \[9x=27|:9\] \[x=3\] ANSWER: Child's ticket costs 3 dollars

OpenStudy (anonymous):

thanks